# LayerPano3D: Layered 3D Panorama for Hyper-Immersive Scene Generation

### v1

@@ Abstract
3D immersive scene generation is a challenging yet critical task in computer vision and graphics. A desired virtual 3D scene should 1) exhibit omnidirectional view consistency, and 2) allow for free exploration in complex scene hierarchies. Existing methods either rely on successive scene expansion via inpainting or employ panorama representation to represent large FOV scene environments. However, the generated scene suffers from semantic drift during expansion and is unable to handle occlusion among scene hierarchies. To tackle these challenges, we introduce LayerPano3D, a novel framework for full-view, explorable panoramic 3D scene generation from a single text prompt. Our key insight is to decompose a reference 2D panorama into multiple layers at different depth levels, where each layer reveals the unseen space from the reference views via diffusion prior. LayerPano3D comprises multiple dedicated designs: 1) we introduce a novel text-guided anchor view synthesis pipeline for high-quality, consistent panorama generation. 2) We pioneer the Layered 3D Panorama as underlying representation to manage complex scene hierarchies and lift it into 3D Gaussians to splat detailed 360-degree omnidirectional scenes with unconstrained viewing paths. Extensive experiments demonstrate that our framework generates state-of-the-art 3D panoramic scene in both full view consistency and immersive exploratory experience. We believe that LayerPano3D holds promise for advancing 3D panoramic scene creation with numerous applications.
Index Terms:
3D Scene Generation, Panorama Generation, Diffusion Models, Neural Rendering
@@
Figure 1: Overview of LayerPano3D. Guided by simple text prompts, LayerPano3D leverages multi-layered 3D panorama to create hyper-immersive panoramic scene with $360^{\circ}\times 180^{\circ}$ coverage, enabling free 3D exploration among complex scene hierarchies.
1 Introduction
@@
The development of spatial computing, including virtual and mixed reality systems, greatly enhances user engagement across various applications, and drives demand for explorable, high-quality 3D environments. We contend that a desired virtual 3D scene should 1) exhibit high-quality and consistency in appearance and geometry across the full $360^{\circ}\times 180^{\circ}$ view; 2) allow for free exploration among complex scene hierarchies with clear parallax. In recent years, many approaches in 3D scene generation [1, 2, 3] were proposed to address these needs.
@@
One branch of works [4, 5, 6, 7, 8, 9] seeks to create extensive scenes by leveraging a “navigate-and-imagine” strategy, which successively applies novel-view rendering and outpaints unseen areas to expand the scene. However, this type of approaches suffer from the semantic drift issue: long sequential scene expansion easily produces incoherent results as the out-paint artifacts accumulate through iterations, hampering the global consistency and harmony of the generated scene.
@@
Another branch of methods [10, 11, 12, 13, 14, 15] employs Equirectangular Panorama to represent $360^{\circ}$, large field of view (FOV) environments in 2D. However, the absence of large-scale panoramic datasets hinders the capability of panorama generation systems, resulting in low-resolution images with simple structures and sparse assets. Moreover, 2D panorama [10, 11, 13] does not allow for free scene exploration. Even when lifted to a panoramic scene [16], the simple spherical structure fails to provide complex scene hierarchies with clear parallax, leading to occluded spaces that cause blurry renderings, ambiguity, and gaps in the generated 3D panorama. Some methods [17] typically use inpainting-based disocclusion strategy to fill in the unseen spaces, but they require specific, predefined rendering paths tailored for each scene, limiting the potential for free exploration.
@@
To this end, we present LayerPano3D, a novel framework that leverages Multi-Layered 3D Panorama for full-view consistent and free exploratory scene generation from text prompts. The main idea is to create a Layered 3D Panorama by first generating a reference panorama and treating it as a multi-layered composition, where each layer depicts scene content at a specific depth level. In this regard, it allows us to create complex scene hierarchies by placing occluded assets in different depth layers at full appearance.
@@
Our contributions are two-fold. First, to generate high-quality and coherent $360^{\circ}\times 180^{\circ}$ panoramas, we propose a novel text-guided anchor view synthesis pipeline. By finetuning a T2I model [18] to generate 4 orthogonal perspective views as anchors, we prevent semantic drifts during panorama generation, while ensuring a consistent horizon level across all views. Furthermore, the anchor views enrich the panorama by incorporating complex structures and detailed features derived from large-scale, pre-trained perspective image generators. Second, we introduce the Layered 3D Panorama representation as a general solution to handle occlusion for different types of scenes with complex scene hierarchies, and lift it to 3D Gaussians [19] to enable free 3D exploration. By leveraging pre-trained panoptic segmentation prior and K-Means clustering, we streamline an automatic layer construction pipeline to decompose the reference panorama into different depth layers. The unseen space at each layer is synthesized with a finetuned panorama inpainter [11].
@@
Extensive experiments demonstrate the effectiveness of LayerPano3D in generating hyper-immersive layered panoramic scene from a single text prompt. LayerPano3D surpasses state-of-the-art methods in creating coherent, plausible, text-aligned 2D panorama and full-view consistent, explorable 3D panoramic environments. Furthermore, our framework streamlines an automatic pipeline without any scene-specific navigation paths, providing more user-friendly interface for non-experts. We believe that LayerPano3D effectively enhances the accessibility of full-view, explorable AIGC 3D environments for real-world applications.
2 Related Works
In recent years, advancements in AI, particularly deep learning, along with increased computational power and large datasets, have driven rapid development in 2D and 3D generation  [20]. These technologies have revolutionized digital content creation in fields like entertainment, design, and virtual reality.
2.1 3D Representation
Early neural rendering methods [21, 22, 23, 24] typically incorporate a multi-layer perceptron (MLP) to model 3D scenes as continuous functions. NeRF [25], being the representative MLP-based method, achieves state-of-the-art novel view synthesis performance due to the volume rendering and inductive bias nature of MLPs. However, as the MLP needs to evaluate on large amount of points along every camera ray, the rendering process becomes extremely exhaustive. Subsequent methods [26, 27, 28, 29, 19] improve on faster training and rendering by employing sparse, advanced scene representation. Among these, 3D Gaussian Splatting (3DGS) [19] has recently emerged as a rising star with transformative real-time rendering and high-definition resolutions. It explicitly parameterizes the 3D scene with 3D Gaussians, where each Gaussian is optimized to represent a volume and projected to 2D for rasterization. The discrete nature of 3D Gaussians enables flexible scene editing and composition for 3DGS. Therefore, our approach employs 3DGS as scene representation to facilitate multi-layer scene creation.
2.2 3D Scene Generation
@@ 2 Related Works
Due to the recent success of diffusion models, 3D scene generation has also achieved some development. Scenescape [7] and DiffDreamer [30], for example, explore perpetual view generation through the incremental construction of 3D scenes. One major branch of work employ step-by-step inpainting from pre-defined trajectories. Text2Room [6] creates room-scale 3D scenes based on text prompt, utilizing textured 3D meshes for scene representation. Similarly, LucidDreamer [4] and WonderJourney [5] can generate domain-free 3D Gaussian splatting scenes from iterative inpainting. However, this line of work often suffer from the semantic drift issue, resulting in unrealistic scene from artifact accumulation and inconsistent semantics. While some other approaches [31, 3, 32] endeavor to integrate objects with environments, they yield relatively low quality of comprehensive scene generation. Recently, our concurrent works, DreamScene360 [16] and HoloDreamer [17] also employ panorama as prior to construct panoramic scenes. However, they only achieve the $360^{\circ}\times 180^{\circ}$ field of view at a fixed viewpoint based on a single panorama of low-quality and simple structure, and do not support free roaming within the scene. In contrast, our framework leverages Multi-Layered 3D Panorama representation to construct high-quality, fully enclosed scenes that enable unconstrained navigation paths in 3D scene.
2.3 Panorama Generation
@@
Panorama generation methods are often based on GANs or diffusion models. Early in this field, with the different forms of deep generative neural networks, GAN-based panorama generation methods explore many paths to improve quality and diversity. Among them, Text2Light [15] focuses on HDR panoramic images by employing a text-conditioned global sampler alongside a structure-aware local sampler. However, training GANs is challenging and they encounter the issue of mode collapse. Recently, some studies have utilized diffusion models to generate panoramas. MVDiffusion [10] generates eight perspective views with multi-branch UNet but the resulting closed-loop panorama only captures the $360^{\circ}\times 90^{\circ}$ FOV. The image generated from MultiDiffusion [33] and Syncdiffusion [34] is more like a long-range image with wide horizontal angle as they do not integrate camera projection models. PanoDiff [35] can generate $360^{\circ}$ panorama from one or more unregistered Narrow Field-of-View (NFoV) images with pose estimation and controlling partial FOV LDM, while the quality and diversity of results are limited by the scarcity of panoramic image training data like most other methods [14, 36, 37]. In contrast, our model can generate Multi-Layered 3D Panorama for immersive, high-quality, and coherent scene generation from text prompts.
3 Method
@@
Figure 2: Pipeline Overview of LayerPano3D. Our framework consists of three stages, namely reference panorama generation, multi-layer panorama construction and panoramic 3D scene optimization.  LayerPano3D streamlines an automatic generation pipeline without any manual efforts to design scene-specific navigation paths for expansion or completion.
@@
The goal of our work is to create a panoramic scene guided by text prompts, that encompasses a complete $360^{\circ}\times 180^{\circ}$ field of view from various viewpoints within an extensive range in the scene, while allowing for unconstrained trajectory for immersive exploration. LayerPano3D consists three stages. In Stage I (Sec. 3.1), we propose a text-guided anchor view synthesis pipeline paired with panoramic outpainting to generate high-quality, consistent panorama as reference. In Stage II (Sec. 3.2), with the reference panorama, we construct our Layered 3D Panorama representation by iterative layer decomposition, completion and alignment process. In Stage III (Sec. 3.3), the Layered 3D Panorama is lifted to 3D Gaussians in a cascaded manner to enable free 3D exploration.
3.1 Reference Panorama Generation
Text-Guided Anchor View Synthesis. We start by generating four orthogonal anchor views to establish the fundamental geometric structure and visual appearance. This is achieved by fine-tuning a pre-trained perspective T2I diffusion model.
For data preparation, we construct a high-quality RGB panorama dataset consisting of 911 outdoor images sourced from the web and 9684 images from Matterport3D [38]. For each panorama, we use llava-v1.5-7b [39] to generate captions as text prompts and project four orthogonal views without overlap at a fixed FOV ($60^{\circ}$), elevation ($0^{\circ}$) and four azimuths ($0^{\circ},90^{\circ},180^{\circ},270^{\circ}$). This setting can effectively avoid perspective distortion and fisheye effect, and allow further refinement of consistency in the subsequent synthesis stages. Additionally, we arrange these four views into a $2\times 2$ grid image in a fixed order for the ease of training and draw a white cross-line with a certain pixel size between the grids as a split.
Compared with object-level multi-view generation, it is harder for scene-level generation to diffuse more views with high quality and consistency in a single pass. Multi-view cross-attention-based methods [10, 11] learn high-level semantics at the expense of overfitting to monotonous patterns due to the scarcity of panorama data. Other approaches [1] require huge computation and time in both training and inference stages and generate relatively simple scenes with limited assets. Therefore, inspired by [40], we directly fine-tune Stable-Diffusion-XL [18] to efficiently generate high-quality anchor views as image grids for subsequent panorama generation, free of complex regularization techniques. This text-guided anchor view synthesis pipeline can scale positively with the potency of the underlying T2I model and unlock the capability of the base model without additional knowledge through lightweight fine-tuning. Furthermore, unlike the clean background in the object-level generation, scene backgrounds are often diverse and complex. Thus, instead of simply initializing each sub-grid with a 2D Gaussian blob as is commonly done in object-based multi-view approaches, we initialize the image grids with random colored noise and draw a white cross-line between the grids during the inference process, mimicking the training data.
Equirectangular Panorama Synthesis. Equirectangular Projection (ERP) is a method that maps a 3D sphere onto a 2D map. In ERP, the lines of longitude are mapped to vertical lines with constant spacing, and the lines of latitude are mapped to horizontal lines with constant spacing. This mapping relationship is neither area-preserving nor angle-preserving and the mapping of latitude leads to strong distortions in the polar regions. To obtain a detailed, consistent and plausible panorama, we first generate a $360^{\circ}\times 90^{\circ}$ FOV panorama to obtain the principle structure and content, then successively extend it to $360^{\circ}\times 180^{\circ}$ FOV panorama.
With the text-guided anchor view synthesis, we have the anchor perspective images $[I_{anc}^{i}]_{i=1}^{4}\in\mathbb{R}^{h\times w\times 3}$ of a scene captured by four orthogonal camera in the horizontal plane. Based on the mapping rule and $60^{\circ}$ FOV, these anchor views can be directly projected to form the incomplete $360^{\circ}\times 60^{\circ}$ FOV panorama without overlap. Here, we assume that the elevation angle of the anchors are $\phi_{anc}=0^{\circ}$ with azimuth angles $\{\theta_{anc}^{i}\}=\{0^{\circ},90^{\circ},180^{\circ},270^{\circ}\}$. Then, we set a viewpoint sequence with $\{(\theta_{j},\phi_{j})\}_{j=1}^{n}$ that allows for a 360-degree rotation around the scene in $90^{\circ}$ FOV to project the perspective views $\hat{I_{j}}$ and masks $M_{j}$:
$[\hat{I_{j}},M_{j}]=\mathcal{ERP}^{-1}(Pano,[\theta_{j},\phi_{j},FOV]),$ (1)
where the $\mathcal{ERP}^{-1}$ is the inverse ERP from panorama to perspective views in 3D. We further apply a text-guided inpainter $\mathcal{F}_{inpaint}$ [41] to synthesize the missing pixels in the masked regions $M_{j}$ of the incomplete views $\hat{I_{j}}$. Note that, $\hat{I_{j}}$ is additionally pre-processed with LaMa [42] before inpainting to clean up the masked pixel distribution with contextual knowledge:
$I_{j}=\mathcal{F}_{inpaint}(LaMa(\hat{I_{j}},M_{j})).$ (2)
Through iterative projection and inpainting, the $Pano_{360\times 60}$ is extended to $Pano_{360\times 90}$.
The extension from $Pano_{360\times 90}$ to $Pano_{360\times 180}$ is more difficult, as the polar regions exhibit most distorted patterns that cannot simply be synthesized through perspective image inpainting. Therefore, we fine-tune [43] on panoramic image-caption pairs for polar outpainting. Another challenge is to enforce continuity between the leftmost and rightmost sides of the synthesized image during outpainting, as it is a crucial characteristic [44] of 360 panoramas. Following  [14, 33], we adopt a circular blending strategy in both denoising and VAE decoding stage, as illustrated in Figure 3. A circular closed loop is created by extending the rightmost side of the panorama with the pixels in the leftmost parts. Per-pixel adaptive weights $W_{k}$ are utilized to blend the denoised patches mapped from extensions and from the leftmost parts to generate the leftmost panorama. Similarly, in the VAE tiled decoding stage, to avoid tiling artifacts, we ensure that each tile overlaps and blends together to form a smooth output due to each tile using a different decoder. Therefore, each pixel of the panorama is computed as a weighted average of all its mapped patch updates $Z_{k,t}$ at each denoising step.
$\Psi(\mathcal{P}_{t})=\sum_{k}\frac{\mathcal{F}_{map,k}^{-1}(W_{k})}{\sum_{m}\mathcal{F}_{map,m}^{-1}(W_{m})}\odot\mathcal{F}_{map,k}^{-1}(\Phi(Z_{k,t})),$ (3)
where $\mathcal{F}_{map,k}^{-1}$ is image space mapping from patch $Z_{k}$ to its corresponding latent in $\mathcal{P}$; $\Psi$ and $\Phi$ denote the diffusion process at panorama-level and patch-level. Finally, we can get a high-quality result as reference for the subsequent layered panorama construction.
Figure 3: The circular blending Strategy. During the inference process, to synthesize a $360^{\circ}$ seamless panorama, adaptive weights are utilized to blend the leftmost part of the latent with the rightmost part in denoising and VAE decoding stage.
3.2 Multi-Layer Panorama Construction

### v2

@@ Abstract.
3D immersive scene generation is a challenging yet critical task in computer vision and graphics. A desired virtual 3D scene should 1) exhibit omnidirectional view consistency, and 2) allow for large-range exploration in complex scene hierarchies. Existing methods either rely on successive scene expansion via inpainting or employ panorama representation to represent large FOV scene environments. However, the generated scene suffers from semantic drift during expansion and is unable to handle occlusion among scene hierarchies. To tackle these challenges, we introduce LayerPano3D, a novel framework for full-view, explorable panoramic 3D scene generation from a single text prompt. Our key insight is to decompose a reference 2D panorama into multiple layers at different depth levels, where each layer reveals the unseen space from the reference views via diffusion prior. LayerPano3D comprises multiple dedicated designs: 1) We introduce a new panorama dataset Upright360 , comprising 9k high-quality and upright panorama images, and finetune the advanced Flux model on Upright360 for high-quality, upright and consistent panorama generation related tasks. 2) We pioneer the Layered 3D Panorama as underlying representation to manage complex scene hierarchies and lift it into 3D Gaussians to splat detailed 360-degree omnidirectional scenes with unconstrained viewing paths. Extensive experiments demonstrate that our framework generates state-of-the-art 3D panoramic scene in both full view consistency and immersive exploratory experience. We believe that LayerPano3D holds promise for advancing 3D panoramic scene creation with numerous applications. More examples please visit our project page: ys-imtech.github.io/projects/LayerPano3D/
journalyear: 2025
@@
Figure 1. Overview of LayerPano3D. Guided by simple text prompts, LayerPano3D leverages multi-layered 3D panorama to create hyper-immersive panoramic scene with $360^{\circ}\times 180^{\circ}$ coverage, enabling free 3D exploration among complex scene hierarchies.
1. Introduction
@@
The development of spatial computing, including virtual and mixed reality systems, greatly enhances user engagement across various applications, and drives demand for explorable, high-quality 3D environments. We contend that a desired virtual 3D scene should 1) exhibit high-quality and consistency in appearance and geometry across the full $360^{\circ}\times 180^{\circ}$ view; 2) allow for exploration among complex scene hierarchies with clear parallax. In recent years, many approaches in 3D scene generation (Gao et al., 2024; Li et al., 2024; Zhang et al., 2023b) were proposed to address these needs.
@@
One branch of works (Chung et al., 2023; Yu et al., 2023; Höllein et al., 2023; Fridman et al., 2023; Ouyang et al., 2023) seeks to create extensive scenes by leveraging a “navigate-and-imagine” strategy, which successively applies novel-view rendering and outpaints unseen areas to expand the scene. However, this type of approaches suffer from the semantic drift issue: long sequential scene expansion easily produces incoherent results as the out-paint artifacts accumulate through iterations, hampering the global consistency and harmony of the generated scene.
@@
Another branch of methods (Tang et al., 2023b; Zhang et al., 2024; Wang et al., 2022, 2023b; Chen et al., 2022) employs Equirectangular Panorama to represent $360^{\circ}$, large field of view (FOV) environments in 2D. However, the absence of large-scale panoramic datasets hinders the capability of panorama generation systems, resulting in low-resolution images with simple structures and sparse assets. Moreover, 2D panorama (Tang et al., 2023b; Zhang et al., 2024; Wang et al., 2022) does not allow for flexible scene exploration. Even when lifted to a panoramic scene (Zhou et al., 2024b), the simple spherical structure fails to provide complex scene hierarchies with clear parallax, leading to occluded spaces that cause blurry renderings, ambiguity, and gaps in the generated 3D panorama. Some methods (Zhou et al., 2024a) typically use inpainting-based disocclusion strategy to fill in the unseen spaces, but they require specific, predefined rendering paths tailored for each scene, limiting the potential for flexible exploration.
@@
To this end, we present LayerPano3D, a novel framework that leverages Multi-Layered 3D Panorama for explorable, full-view consistent scene generation from text prompts. The main idea is to create a Layered 3D Panorama by first generating a reference panorama and treating it as a multi-layered composition, where each layer depicts scene content at a specific depth level. In this regard, it allows us to create complex scene hierarchies by placing occluded assets in different depth layers at full appearance.
Our contributions are two-fold. First, to generate high-quality and coherent $360^{\circ}\times 180^{\circ}$ panoramas, we curate a new dataset, namely Upright360 , consisting of 9k high-quality, upright panorama images, and finetune the advanced Flux (Labs, 2023) model with panorama LoRA on it for panorama generation and inpainting. This feed-forward pipeline prevents semantic drifts during panorama generation, while ensuring a consistent horizon level across all views.
@@
Second, we introduce the Layered 3D Panorama representation as a general solution to handle occlusion for different types of scenes with complex scene hierarchies, and lift it to 3D Gaussians (Kerbl et al., 2023) to enable large-range 3D exploration. By leveraging pre-trained panoptic segmentation prior and K-Means clustering, we streamline an automatic layer construction pipeline to decompose the reference panorama into different depth layers. The unseen space at each layer is synthesized with the Flux-based inpainting pipeline.
@@
Extensive experiments demonstrate the effectiveness of LayerPano3D in generating hyper-immersive layered panoramic scene from a single text prompt. LayerPano3D surpasses state-of-the-art methods in creating coherent, plausible, text-aligned 2D panorama and full-view consistent, explorable 3D panoramic environments. Furthermore, our framework does not require any scene-specific navigation paths, providing more user-friendly interface for non-experts. We believe that LayerPano3D effectively enhances the accessibility of full-view, explorable AIGC 3D environments for real-world applications.
2. Related Works
2.1. 3D Scene Generation
@@ 2. Related Works
Due to the recent success of diffusion models (Tang et al., 2023a; Poole et al., 2022), 3D scene generation has also achieved some development. Scenescape (Fridman et al., 2023) and DiffDreamer (Cai et al., 2023), for example, explore perpetual view generation through the incremental construction of 3D scenes. One major branch of work employ step-by-step inpainting from pre-defined trajectories. Text2Room (Höllein et al., 2023) creates room-scale 3D scenes based on text prompt, utilizing textured 3D meshes for scene representation. Similarly, LucidDreamer (Chung et al., 2023) and WonderJourney (Yu et al., 2023) can generate domain-free 3D Gaussian splatting scenes from iterative inpainting. However, this line of work often suffer from the semantic drift issue, resulting in unrealistic scene from artifact accumulation and inconsistent semantics. While some other approaches (Cohen-Bar et al., 2023; Zhang et al., 2023b; Vilesov et al., 2023) endeavor to integrate objects with environments, they yield relatively low quality of comprehensive scene generation. Recently, our concurrent works, DreamScene360 (Zhou et al., 2024b) and HoloDreamer (Zhou et al., 2024a) also employ panorama as prior to construct panoramic scenes. However, they only achieve the $360^{\circ}\times 180^{\circ}$ field of view at a fixed viewpoint based on a single panorama of low-quality and simple structure, and do not support free roaming within the scene. In contrast, our framework leverages Multi-Layered 3D Panorama representation to construct high-quality, fully enclosed scenes that enable larger-range exploration paths in 3D scene.
2.2. Panorama Generation
@@
Panorama generation methods are often based on GANs or diffusion models. Early in this field, with the different forms of deep generative neural networks, GAN-based panorama generation methods explore many paths to improve quality and diversity. Among them, Text2Light (Chen et al., 2022) focuses on HDR panoramic images by employing a text-conditioned global sampler alongside a structure-aware local sampler. However, training GANs is challenging and they encounter the issue of mode collapse. Recently, some studies have utilized diffusion models to generate panoramas. MVDiffusion (Tang et al., 2023b) generates eight perspective views with multi-branch UNet but the resulting closed-loop panorama only captures the $360^{\circ}\times 90^{\circ}$ FOV. The image generated from MultiDiffusion (Bar-Tal et al., 2023) and Syncdiffusion (Lee et al., 2023) is more like a long-range image with wide horizontal angle as they do not integrate camera projection models. PanoDiff (Wang et al., 2023a) can generate $360^{\circ}$ panorama from one or more unregistered Narrow Field-of-View (NFoV) images with pose estimation and controlling partial FOV LDM, while the quality and diversity of results are limited by the scarcity of panoramic image training data like most other methods (Wang et al., 2023b; Li and Bansal, 2023; Wu et al., 2024). In contrast, our model can generate Multi-Layered 3D Panorama for immersive, high-quality, and coherent scene generation from text prompts.
3. Method
@@
Figure 2. Pipeline Overview of LayerPano3D. Our framework consists of two stages, namely multi-layer panorama construction and panoramic 3D scene optimization.  LayerPano3D streamlines an automatic generation pipeline without any manual efforts to design scene-specific navigation paths for expansion or completion.
@@
The goal of our work is to create a panoramic scene guided by text prompts. This generated scene encompasses a complete $360^{\circ}\times 180^{\circ}$ field of view from various viewpoints within an extensive range in the scene, while allowing for immersive exploration along complex trajectories. LayerPano3D consists two stages. In Stage I (Sec. 3.1), we first generate the reference panorama from text prompt by finetuning flux (Labs, 2023) with panorama LoRA in our Upright360 dataset. With the reference panorama, we construct our Layered 3D Panorama representation by iterative layer decomposition, completion and alignment process. In Stage II (Sec. 3.2), the Layered 3D Panorama is lifted to 3D Gaussians in a cascaded manner to enable large-range 3D exploration.
3.1. Multi-Layer Panorama Generation